OpenStudy (anonymous):

f(x) = x^5 + 3x^3 + 2x - 1; P(5,1) a) Use f' to prove that f has an inverse function b) Find the slope of the tangent line at the point P on the graph of f-1

OpenStudy (bahrom7893):

http://en.wikipedia.org/wiki/Inverse_functions_and_differentiation According to that \[\frac{dx}{dy}*\frac{dy}{dx}=1\]Find both of those.

OpenStudy (bahrom7893):

\[y = x^5 + 3x^3 + 2x - 1\]\[\frac{dy}{dx}=5x^4+9x^2+2\]

OpenStudy (bahrom7893):

Now use implicit differentiation again to find \[\frac{dx}{dy}\]

OpenStudy (anonymous):

ok, so Dy/dx is increasing, so it's on-to-one, and therefore a function! but I don't even know how to find the slope per the question...I appreciate your time here!

OpenStudy (bahrom7893):

You need to find \[\frac{dx}{dy}\]which should be easy to find via implicit differentiation.. For some reason I'm blanking on that right now. Once you find dx/dy, you can just plug in your point and get the slope.

OpenStudy (anonymous):

ok. thanks, I think I can take it from here.

OpenStudy (anonymous):

Implicit differentiation is not one of my strong point,,guess this is a good time to review!